# Continuous-Time Quantum Markov Chains And Discretizations Of $p$-Adic Schrödinger Equations: Comparisons And Simulations

### v1

@@ 7. Numerical simulations
$p_{r,v}(t)=p^{l}{\displaystyle\int\limits_{B_{-l}(v)}}Z_{0}\left(x,t\right)\ast\Omega\left(p^{l}\left|x-r\right|_{p}\right)dx=\left\langle\Omega\left(p^{l}\left|x-v\right|_{p}\right),Z_{0}\left(x,t\right)\ast\Omega\left(p^{l}\left|x-r\right|_{p}\right)\right\rangle,$
@@
$Z_{0}\left(x,t\right)\ast\Omega\left(p^{l}\left|x-r\right|_{p}\right)=e^{-t\boldsymbol{J}}\Omega\left(p^{l}\left|x-r\right|_{p}\right)=e^{t\boldsymbol{J}^{\left(l\right)}}e_{r},$
@@
(7.1) $p_{r,v}(t)=\left\langle e_{v},e^{t\boldsymbol{J}^{\left(l\right)}}e_{r}\right\rangle_{l}.$
@@ 7.1. Heat and Schrödinger equation attached to Bessel potentials
Then $\left\|J_{\alpha}\right\|_{1}=1$, $J_{\alpha}\left(|\xi|_{p}\right)\geq 0$, for $x\neq 0$. Furthermore, $\widehat{J}_{\alpha}\left(|\xi|_{p}\right)=\left(\max\left\{1,\left|\xi\right|_{p}\right\}\right)^{-\alpha}$, see [41, Lemma 5.2].
@@
$i\frac{\partial}{\partial t}\Psi\left(x,t\right)=-\left\{J_{\alpha}\left(\left|x\right|_{p}\right)\ast\Psi\left(x,t\right)-\Psi\left(x,t\right)\right\}\text{, }x\in\mathbb{Z}_{p},t\geq 0.$
@@ 7.2. Numerical simulation 1
$p_{I,J}(t)=\left\langle e_{J},e^{t\boldsymbol{J}^{\left(l\right)}\left(\alpha\right)}e_{I}\right\rangle_{l}$
@@
$\pi_{I,J}\left(t\right)=\left|\left\langle e_{I}\right|e^{it\boldsymbol{J}^{\left(l\right)}\left(\alpha\right)}\left|e_{J}\right\rangle_{l}\right|^{2}\text{, for }J,I\in G_{l},$
@@ 7.3. Numerical simulation 2
$p_{r,v}(t)=\left\langle e_{v},e^{t\boldsymbol{J}^{\left(l\right)}\left(1\right)}e_{r}\right\rangle_{l}.$
@@
$\pi_{I,J}\left(t\right)=\left|\left\langle e_{I}\right|e^{it\boldsymbol{J}^{\left(l\right)}\left(1\right)}\left|e_{J}\right\rangle_{l}\right|^{2}\text{, for }J,I\in G_{l}.$
@@ 9.1. Preliminary results
$Z_{0}\left(x,t\right):=\left\{\begin{array}[c]{ll}\mathcal{F}_{\xi\rightarrow x}^{-1}\left(e^{-t\left(1-\widehat{J}(\left|\xi\right|_{p})\right)}\right)\in\mathcal{D}^{\prime}\left(\mathbb{Z}_{p}\right),&\text{if }t>0.\\ &\\ \delta\left(x\right)&\text{if }t=0,\end{array}\right$.
@@
(9.2) $Z_{0}\left(x,t\right)$ $=\Omega\left(\left|x\right|_{p}\right)+{\displaystyle\sum\limits_{j=1}^{\infty}}e^{-t\left(1-\widehat{J}\left(p^{j}\right)\right)}\left\{p^{j}\Omega\left(p^{j}\left|x\right|_{p}\right)-p^{\left(j-1\right)}\Omega\left(p^{\left(j-1\right)}\left|x\right|_{p}\right)\right\}$
$=\Omega\left(\left|x\right|_{p}\right)-e^{-t\left(1-\widehat{J}\left(p\right)\right)}1_{\mathbb{Z}_{p}^{\times}}\left(x\right)+\Omega\left(p\left|x\right|_{p}\right){\displaystyle\sum\limits_{j=1}^{\infty}}p^{j}e^{-t\left(1-\widehat{J}\left(p^{j}\right)\right)},$
@@
$\left(1-p^{-1}\right){\displaystyle\sum\limits_{j=1}^{\infty}}p^{j}e^{-t\left(1-\widehat{J}\left(p^{j}\right)\right)}$
@@
$e^{-t\left(1-\widehat{J}(\left|\xi\right|_{p})\right)}$ $=\Omega\left(\left|\xi\right|_{p}\right)+{\displaystyle\sum\limits_{j=1}^{\infty}}e^{-t\left(1-\widehat{J}(p^{j})\right)}1_{S_{j}}\left(\xi\right)$
$=\Omega\left(\left|\xi\right|_{p}\right)+{\displaystyle\sum\limits_{j=1}^{\infty}}e^{-t\left(1-\widehat{J}(p^{j})\right)}\left\{\Omega\left(p^{-j}\left|x\right|_{p}\right)-\Omega\left(p^{-\left(j-1\right)}\left|x\right|_{p}\right)\right\}.$
@@
$\Omega\left(\left|\xi\right|_{p}\right)+{\displaystyle\sum\limits_{j=1}^{m}}e^{-t\left(1-\widehat{J}(p^{j})\right)}\left\{\Omega\left(p^{-j}\left|x\right|_{p}\right)-\Omega\left(p^{-\left(j-1\right)}\left|x\right|_{p}\right)\right\}\rightarrow e^{-t\left(1-\widehat{J}(\left|\xi\right|_{p})\right)}$
@@
$\mathcal{F}_{\xi\rightarrow x}^{-1}\left(\Omega\left(\left|\xi\right|_{p}\right)\right)+{\displaystyle\sum\limits_{j=1}^{m}}e^{-t\left(1-\widehat{J}(p^{j})\right)}\mathcal{F}_{\xi\rightarrow x}^{-1}\left\{\Omega\left(p^{-j}\left|x\right|_{p}\right)-\Omega\left(p^{-\left(j-1\right)}\left|x\right|_{p}\right)\right\}\\ \rightarrow\mathcal{F}_{\xi\rightarrow x}^{-1}\left(e^{-t\left(1-\widehat{J}(\left|\xi\right|_{p})\right)}\right)$
@@
$\mathcal{F}_{\xi\rightarrow x}^{-1}\left(\Omega\left(\left|\xi\right|_{p}\right)\right)+{\displaystyle\sum\limits_{j=1}^{m}}e^{-t\left(1-\widehat{J}(p^{j})\right)}\mathcal{F}_{\xi\rightarrow x}^{-1}\left\{\Omega\left(p^{-j}\left|x\right|_{p}\right)-\Omega\left(p^{-\left(j-1\right)}\left|x\right|_{p}\right)\right\}$
$=\Omega\left(\left|x\right|_{p}\right)+{\displaystyle\sum\limits_{j=1}^{m}}e^{-t\left(1-\widehat{J}\left(p^{j}\right)\right)}\left\{p^{j}\Omega\left(p^{j}\left|x\right|_{p}\right)-p^{\left(j-1\right)}\Omega\left(p^{\left(j-1\right)}\left|x\right|_{p}\right)\right\},$
@@
$\boldsymbol{J}\left(Z_{0}\left(x,t\right)\right):=\mathcal{F}_{\xi\rightarrow x}^{-1}\left(\left\{\widehat{J}(\left|\xi\right|_{p})-1\right\}e^{-t\left(1-\widehat{J}(\left|\xi\right|_{p})\right)}\right)\in\mathcal{D}^{\prime}\left(\mathbb{Z}_{p}\right),$
@@
$\boldsymbol{J}\left(Z_{0}\left(x,t\right)\right)={\displaystyle\sum\limits_{j=1}^{\infty}}e^{-t\left(1-\widehat{J}\left(p^{j}\right)\right)}\left(\widehat{J}\left(p^{j}\right)-1\right)\left\{p^{j}\Omega\left(p^{j}\left|x\right|_{p}\right)-p^{\left(j-1\right)}\Omega\left(p^{\left(j-1\right)}\left|x\right|_{p}\right)\right\},$
@@
$\left(Z_{0}(x,t),\Omega\left(\left|x\right|_{p}\right)\right)$ $=\left(\mathcal{F}_{\xi\rightarrow x}^{-1}\left(e^{-t\left(1-\widehat{J}(\left|\xi\right|_{p})\right)}\right),\Omega\left(\left|x\right|_{p}\right)\right)=\left(e^{-t\left(1-\widehat{J}(\left|\xi\right|_{p})\right)},\Omega\left(\left|\xi\right|_{p}\right)\right)$
$={\displaystyle\int\limits_{\mathbb{Z}_{p}}}e^{-t\left(1-\widehat{J}(\left|\xi\right|_{p})\right)}d\xi=e^{-t\left(1-\widehat{J}(0)\right)}{\displaystyle\int\limits_{\mathbb{Z}_{p}}}d\xi=1.$
@@
$Z_{0}^{\left(m\right)}\left(x,t\right)=\Omega\left(\left|x\right|_{p}\right)+{\displaystyle\sum\limits_{j=1}^{m}}e^{-t\left(1-\widehat{J}\left(p^{j}\right)\right)}\left\{p^{j}\Omega\left(p^{j}\left|x\right|_{p}\right)-p^{\left(j-1\right)}\Omega\left(p^{\left(j-1\right)}\left|x\right|_{p}\right)\right\}.$
@@ 10. Appendix C: Continuous-time Markov chains
$\leq e^{-t\left(1-\widehat{J}(0)\right)}\Omega\left(\left|x\right|_{p}\right)=\Omega\left(\left|x\right|_{p}\right)\text{, }$
@@
${\displaystyle\sum\limits_{r\in\mathbb{J}}}p_{r,v}(t)=\\ \frac{1}{\mu_{\text{Haar}}\left(\mathcal{K}_{v}\right)}\lim_{n\rightarrow\infty}{\displaystyle\sum\limits_{r\in\mathbb{J}_{n}}}\text{ }{\displaystyle\int\limits_{\mathcal{K}_{v}}}Z_{0}\left(x,t\right)\ast 1_{\mathcal{K}_{r}}\left(x\right)dx=\frac{1}{\mu_{\text{Haar}}\left(\mathcal{K}_{v}\right)}\lim_{n\rightarrow\infty}{\displaystyle\int\limits_{\mathcal{K}_{v}}}Z_{0}\left(x,t\right)\ast{\displaystyle\sum\limits_{r\in\mathbb{J}_{n}}}1_{\mathcal{K}_{r}}\left(x\right)dx\\ =\frac{1}{\mu_{\text{Haar}}\left(\mathcal{K}_{v}\right)}{\displaystyle\int\limits_{\mathcal{K}_{v}}}Z_{0}\left(x,t\right)\ast\lim_{n\rightarrow\infty}{\displaystyle\sum\limits_{r\in\mathbb{J}_{n}}}1_{\mathcal{K}_{r}}\left(x\right)dx=\frac{1}{\mu_{\text{Haar}}\left(\mathcal{K}_{v}\right)}{\displaystyle\int\limits_{\mathcal{K}_{v}}}Z_{0}\left(x,t\right)\ast\Omega\left(\left|x\right|_{p}\right)dx\\ =\frac{1}{\mu_{\text{Haar}}\left(\mathcal{K}_{v}\right)}{\displaystyle\int\limits_{\mathcal{K}_{v}}}\left\{{\displaystyle\int\limits_{\mathbb{Z}_{p}}}Z_{0}\left(x-y,t\right)dy\right\}dx=\frac{1}{\mu_{\text{Haar}}\left(\mathcal{K}_{v}\right)}{\displaystyle\int\limits_{\mathcal{K}_{v}}}\left\{{\displaystyle\int\limits_{\mathbb{Z}_{p}}}Z_{0}\left(u,t\right)du\right\}dx\\ =\frac{e^{-t\left(1-\widehat{J}(0)\right)}\int_{\mathcal{K}_{v}}dx}{\mu_{\text{Haar}}\left(\mathcal{K}_{v}\right)}=1.$
@@
$\widehat{F}(\xi,t)$ $=\widehat{1}_{\mathcal{K}_{v}}\left(\xi\right)\ast\left\{e^{-t\left(1-\widehat{J}(\left|\xi\right|_{p})\right)}-\Omega\left(\left|\xi\right|_{p}\right)\right\}\widehat{1}_{\mathcal{K}_{r}}\left(\xi\right)$
$=\widehat{1}_{\mathcal{K}_{v}}\left(\xi\right)\ast\left\{\begin{array}[c]{lll}0&\text{if }&\left|\xi\right|_{p}\leq 1\\ &&\\ \widehat{1}_{\mathcal{K}_{r}}\left(\xi\right)e^{-t\left(1-\widehat{J}(\left|\xi\right|_{p})\right)}&\text{if }&\left|\xi\right|_{p}\geq p\end{array}\right$.
$={\displaystyle\int\limits_{\mathbb{Q}_{p}\smallsetminus\mathbb{Z}_{p}}}\widehat{1}_{\mathcal{K}_{v}}\left(\xi-u\right)e^{-t\left(1-\widehat{J}(\left|u\right|_{p})\right)}\widehat{1}_{\mathcal{K}_{r}}\left(u\right)du.$
Now, by the hypothesis $1-\widehat{J}(\left|u\right|_{p})>0$ for $u\in\mathbb{Q}_{p}\smallsetminus\mathbb{Z}_{p}$, $e^{-t\left(1-\widehat{J}(\left|u\right|_{p})\right)}\leq 1$ for $t>0$, and the fact that $\widehat{1}_{\mathcal{K}_{v}}\left(-u\right)\widehat{1}_{\mathcal{K}_{r}}\left(u\right)$ is a test function, we can apply the dominated convergence theorem,
$\lim_{t\rightarrow\infty}\widehat{F}(0,t)$ $=\lim_{t\rightarrow\infty}{\displaystyle\int\limits_{\mathbb{Q}_{p}\smallsetminus\mathbb{Z}_{p}}}\widehat{1}_{\mathcal{K}_{v}}\left(-u\right)e^{-t\left(1-\widehat{J}(\left|u\right|_{p})\right)}\widehat{1}_{\mathcal{K}_{r}}\left(u\right)du$
$={\displaystyle\int\limits_{\mathbb{Q}_{p}\smallsetminus\mathbb{Z}_{p}}}\widehat{1}_{\mathcal{K}_{v}}\left(-u\right)\left\{\lim_{t\rightarrow\infty}e^{-t\left(1-\widehat{J}(\left|u\right|_{p})\right)}\right\}\widehat{1}_{\mathcal{K}_{r}}\left(u\right)du=0.$

### v2

@@ 7. Numerical simulations
$p_{r,v}(t)$ $=$ $p^{l}\int\limits_{B_{-l}(v)}Z_{0}\left(x,t\right)\ast\Omega\left(p^{l}\left|x-r\right|_{p}\right)dx$
$=$ $\left\langle\Omega\left(p^{l}\left|x-v\right|_{p}\right),Z_{0}\left(x,t\right)\ast\Omega\left(p^{l}\left|x-r\right|_{p}\right)\right\rangle,$
@@
$Z_{0}\left(x,t\right)\ast\Omega\left(p^{l}\left|x-r\right|_{p}\right)=\mathrm{e}^{-t\boldsymbol{J}}\Omega\left(p^{l}\left|x-r\right|_{p}\right)=\mathrm{e}^{t\boldsymbol{J}^{\left(l\right)}}e_{r},$
@@
(7.1) $p_{r,v}(t)=\left\langle e_{v},\mathrm{e}^{t\boldsymbol{J}^{\left(l\right)}}e_{r}\right\rangle_{l}.$
@@ 7.1. Heat and Schrödinger equation attached to Bessel potentials
Then $\left\|J_{\alpha}\right\|_{1}=1$, $J_{\alpha}\left(|\xi|_{p}\right)\geq 0$, for $x\neq 0$, and $\widehat{J}_{\alpha}\left(|\xi|_{p}\right)=\left(\max\left\{1,\left|\xi\right|_{p}\right\}\right)^{-\alpha}$, see [41, Lemma 5.2].
@@
$\mathrm{i}\frac{\partial}{\partial t}\Psi\left(x,t\right)=-\left\{J_{\alpha}\left(\left|x\right|_{p}\right)\ast\Psi\left(x,t\right)-\Psi\left(x,t\right)\right\}\text{, }x\in\mathbb{Z}_{p},t\geq 0.$
@@ 7.2. Numerical simulation 1
$p_{I,J}(t)=\left\langle e_{J},\mathrm{e}^{t\boldsymbol{J}^{\left(l\right)}\left(\alpha\right)}e_{I}\right\rangle_{l}$
@@
$\pi_{I,J}\left(t\right)=\left|\left\langle e_{I}\right|\mathrm{e}^{\mathrm{i}t\boldsymbol{J}^{\left(l\right)}\left(\alpha\right)}\left|e_{J}\right\rangle_{l}\right|^{2}\text{, for }J,I\in G_{l},$
@@ 7.3. Numerical simulation 2
$p_{r,v}(t)=\left\langle e_{v},\mathrm{e}^{t\boldsymbol{J}^{\left(l\right)}\left(1\right)}e_{r}\right\rangle_{l}.$
@@
$\pi_{I,J}\left(t\right)=\left|\left\langle e_{I}\right|\mathrm{e}^{\mathrm{i}t\boldsymbol{J}^{\left(l\right)}\left(1\right)}\left|e_{J}\right\rangle_{l}\right|^{2}\text{, for }J,I\in G_{l}.$
@@ 9.1. Preliminary results
$Z_{0}\left(x,t\right):=\left\{\begin{array}[c]{ll}\mathcal{F}_{\xi\rightarrow x}^{-1}\left(\mathrm{e}^{-t\left(1-\widehat{J}(\left|\xi\right|_{p})\right)}\right)\in\mathcal{D}^{\prime}\left(\mathbb{Z}_{p}\right),&\text{if }t>0.\\ &\\ \delta\left(x\right)&\text{if }t=0,\end{array}\right$.
@@
(9.2) $Z_{0}\left(x,t\right)$ $=\Omega\left(\left|x\right|_{p}\right)+{\displaystyle\sum\limits_{j=1}^{\infty}}\mathrm{e}^{-t\left(1-\widehat{J}\left(p^{j}\right)\right)}\left\{p^{j}\Omega\left(p^{j}\left|x\right|_{p}\right)-p^{\left(j-1\right)}\Omega\left(p^{\left(j-1\right)}\left|x\right|_{p}\right)\right\}$
$=\Omega\left(\left|x\right|_{p}\right)-\mathrm{e}^{-t\left(1-\widehat{J}\left(p\right)\right)}1_{\mathbb{Z}_{p}^{\times}}\left(x\right)+\Omega\left(p\left|x\right|_{p}\right){\displaystyle\sum\limits_{j=1}^{\infty}}p^{j}\mathrm{e}^{-t\left(1-\widehat{J}\left(p^{j}\right)\right)},$
@@
$\left(1-p^{-1}\right){\displaystyle\sum\limits_{j=1}^{\infty}}p^{j}\mathrm{e}^{-t\left(1-\widehat{J}\left(p^{j}\right)\right)}$
@@
$\mathrm{e}^{-t\left(1-\widehat{J}(\left|\xi\right|_{p})\right)}$ $=\Omega\left(\left|\xi\right|_{p}\right)+{\displaystyle\sum\limits_{j=1}^{\infty}}\mathrm{e}^{-t\left(1-\widehat{J}(p^{j})\right)}1_{S_{j}}\left(\xi\right)$
$=\Omega\left(\left|\xi\right|_{p}\right)+{\displaystyle\sum\limits_{j=1}^{\infty}}\mathrm{e}^{-t\left(1-\widehat{J}(p^{j})\right)}\left\{\Omega\left(p^{-j}\left|x\right|_{p}\right)-\Omega\left(p^{-\left(j-1\right)}\left|x\right|_{p}\right)\right\}.$
@@
$\Omega\left(\left|\xi\right|_{p}\right)+{\displaystyle\sum\limits_{j=1}^{m}}\mathrm{e}^{-t\left(1-\widehat{J}(p^{j})\right)}\left\{\Omega\left(p^{-j}\left|x\right|_{p}\right)-\Omega\left(p^{-\left(j-1\right)}\left|x\right|_{p}\right)\right\}\rightarrow\mathrm{e}^{-t\left(1-\widehat{J}(\left|\xi\right|_{p})\right)}$
@@
$\mathcal{F}_{\xi\rightarrow x}^{-1}\left(\Omega\left(\left|\xi\right|_{p}\right)\right)+{\displaystyle\sum\limits_{j=1}^{m}}\mathrm{e}^{-t\left(1-\widehat{J}(p^{j})\right)}\mathcal{F}_{\xi\rightarrow x}^{-1}\left\{\Omega\left(p^{-j}\left|x\right|_{p}\right)-\Omega\left(p^{-\left(j-1\right)}\left|x\right|_{p}\right)\right\}\\ \rightarrow\mathcal{F}_{\xi\rightarrow x}^{-1}\left(\mathrm{e}^{-t\left(1-\widehat{J}(\left|\xi\right|_{p})\right)}\right)$
@@
$\mathcal{F}_{\xi\rightarrow x}^{-1}\left(\Omega\left(\left|\xi\right|_{p}\right)\right)+{\displaystyle\sum\limits_{j=1}^{m}}\mathrm{e}^{-t\left(1-\widehat{J}(p^{j})\right)}\mathcal{F}_{\xi\rightarrow x}^{-1}\left\{\Omega\left(p^{-j}\left|x\right|_{p}\right)-\Omega\left(p^{-\left(j-1\right)}\left|x\right|_{p}\right)\right\}$
$=\Omega\left(\left|x\right|_{p}\right)+{\displaystyle\sum\limits_{j=1}^{m}}\mathrm{e}^{-t\left(1-\widehat{J}\left(p^{j}\right)\right)}\left\{p^{j}\Omega\left(p^{j}\left|x\right|_{p}\right)-p^{\left(j-1\right)}\Omega\left(p^{\left(j-1\right)}\left|x\right|_{p}\right)\right\},$
@@
$\boldsymbol{J}\left(Z_{0}\left(x,t\right)\right):=\mathcal{F}_{\xi\rightarrow x}^{-1}\left(\left\{\widehat{J}(\left|\xi\right|_{p})-1\right\}\mathrm{e}^{-t\left(1-\widehat{J}(\left|\xi\right|_{p})\right)}\right)\in\mathcal{D}^{\prime}\left(\mathbb{Z}_{p}\right),$
@@
$\boldsymbol{J}\left(Z_{0}\left(x,t\right)\right)={\displaystyle\sum\limits_{j=1}^{\infty}}\mathrm{e}^{-t\left(1-\widehat{J}\left(p^{j}\right)\right)}\left(\widehat{J}\left(p^{j}\right)-1\right)\left\{p^{j}\Omega\left(p^{j}\left|x\right|_{p}\right)-p^{\left(j-1\right)}\Omega\left(p^{\left(j-1\right)}\left|x\right|_{p}\right)\right\},$
@@
$\left(Z_{0}(x,t),\Omega\left(\left|x\right|_{p}\right)\right)$ $=\left(\mathcal{F}_{\xi\rightarrow x}^{-1}\left(\mathrm{e}^{-t\left(1-\widehat{J}(\left|\xi\right|_{p})\right)}\right),\Omega\left(\left|x\right|_{p}\right)\right)=\left(\mathrm{e}^{-t\left(1-\widehat{J}(\left|\xi\right|_{p})\right)},\Omega\left(\left|\xi\right|_{p}\right)\right)$
$={\displaystyle\int\limits_{\mathbb{Z}_{p}}}\mathrm{e}^{-t\left(1-\widehat{J}(\left|\xi\right|_{p})\right)}d\xi=\mathrm{e}^{-t\left(1-\widehat{J}(0)\right)}{\displaystyle\int\limits_{\mathbb{Z}_{p}}}d\xi=1.$
@@
$Z_{0}^{\left(m\right)}\left(x,t\right)=\Omega\left(\left|x\right|_{p}\right)+{\displaystyle\sum\limits_{j=1}^{m}}\mathrm{e}^{-t\left(1-\widehat{J}\left(p^{j}\right)\right)}\left\{p^{j}\Omega\left(p^{j}\left|x\right|_{p}\right)-p^{\left(j-1\right)}\Omega\left(p^{\left(j-1\right)}\left|x\right|_{p}\right)\right\}.$
@@ 10. Appendix C: Continuous-time Markov chains
$\leq\mathrm{e}^{-t\left(1-\widehat{J}(0)\right)}\Omega\left(\left|x\right|_{p}\right)=\Omega\left(\left|x\right|_{p}\right)\text{, }$
@@
$\sum\limits_{r\in\mathbb{J}}p_{r,v}(t)=\frac{1}{\mu_{\text{Haar}}\left(\mathcal{K}_{v}\right)}\lim_{n\rightarrow\infty}\sum\limits_{r\in\mathbb{J}_{n}}\text{ }\int\limits_{\mathcal{K}_{v}}Z_{0}\left(x,t\right)\ast 1_{\mathcal{K}_{r}}\left(x\right)dx\\ =\frac{1}{\mu_{\text{Haar}}\left(\mathcal{K}_{v}\right)}\lim_{n\rightarrow\infty}\int\limits_{\mathcal{K}_{v}}Z_{0}\left(x,t\right)\ast\sum\limits_{r\in\mathbb{J}_{n}}1_{\mathcal{K}_{r}}\left(x\right)dx\\ =\frac{1}{\mu_{\text{Haar}}\left(\mathcal{K}_{v}\right)}\int\limits_{\mathcal{K}_{v}}Z_{0}\left(x,t\right)\ast\lim_{n\rightarrow\infty}\sum\limits_{r\in\mathbb{J}_{n}}1_{\mathcal{K}_{r}}\left(x\right)dx=\frac{1}{\mu_{\text{Haar}}\left(\mathcal{K}_{v}\right)}\int\limits_{\mathcal{K}_{v}}Z_{0}\left(x,t\right)\ast\Omega\left(\left|x\right|_{p}\right)dx\\ =\frac{1}{\mu_{\text{Haar}}\left(\mathcal{K}_{v}\right)}\int\limits_{\mathcal{K}_{v}}\left\{\int\limits_{\mathbb{Z}_{p}}Z_{0}\left(x-y,t\right)dy\right\}dx=\frac{1}{\mu_{\text{Haar}}\left(\mathcal{K}_{v}\right)}\int\limits_{\mathcal{K}_{v}}\left\{\int\limits_{\mathbb{Z}_{p}}Z_{0}\left(u,t\right)du\right\}dx\\ =\frac{\mathrm{e}^{-t\left(1-\widehat{J}(0)\right)}\int_{\mathcal{K}_{v}}dx}{\mu_{\text{Haar}}\left(\mathcal{K}_{v}\right)}=1.$
@@
$\widehat{F}(\xi,t)$ $=\widehat{1}_{\mathcal{K}_{v}}\left(\xi\right)\ast\left\{\mathrm{e}^{-t\left(1-\widehat{J}(\left|\xi\right|_{p})\right)}-\Omega\left(\left|\xi\right|_{p}\right)\right\}\widehat{1}_{\mathcal{K}_{r}}\left(\xi\right)$
$=\widehat{1}_{\mathcal{K}_{v}}\left(\xi\right)\ast\left\{\begin{array}[c]{lll}0&\text{if }&\left|\xi\right|_{p}\leq 1\\ &&\\ \widehat{1}_{\mathcal{K}_{r}}\left(\xi\right)\mathrm{e}^{-t\left(1-\widehat{J}(\left|\xi\right|_{p})\right)}&\text{if }&\left|\xi\right|_{p}\geq p\end{array}\right$.
$={\displaystyle\int\limits_{\mathbb{Q}_{p}\smallsetminus\mathbb{Z}_{p}}}\widehat{1}_{\mathcal{K}_{v}}\left(\xi-u\right)\mathrm{e}^{-t\left(1-\widehat{J}(\left|u\right|_{p})\right)}\widehat{1}_{\mathcal{K}_{r}}\left(u\right)du.$
Now, by the hypothesis $1-\widehat{J}(\left|u\right|_{p})>0$ for $u\in\mathbb{Q}_{p}\smallsetminus\mathbb{Z}_{p}$, $\mathrm{e}^{-t\left(1-\widehat{J}(\left|u\right|_{p})\right)}\leq 1$ for $t>0$, and the fact that $\widehat{1}_{\mathcal{K}_{v}}\left(-u\right)\widehat{1}_{\mathcal{K}_{r}}\left(u\right)$ is a test function, we can apply the dominated convergence theorem,
$\lim_{t\rightarrow\infty}\widehat{F}(0,t)$ $=\lim_{t\rightarrow\infty}{\displaystyle\int\limits_{\mathbb{Q}_{p}\smallsetminus\mathbb{Z}_{p}}}\widehat{1}_{\mathcal{K}_{v}}\left(-u\right)\mathrm{e}^{-t\left(1-\widehat{J}(\left|u\right|_{p})\right)}\widehat{1}_{\mathcal{K}_{r}}\left(u\right)du$
$={\displaystyle\int\limits_{\mathbb{Q}_{p}\smallsetminus\mathbb{Z}_{p}}}\widehat{1}_{\mathcal{K}_{v}}\left(-u\right)\left\{\lim_{t\rightarrow\infty}\mathrm{e}^{-t\left(1-\widehat{J}(\left|u\right|_{p})\right)}\right\}\widehat{1}_{\mathcal{K}_{r}}\left(u\right)du=0.$